OpenStudy (anonymous):

John was graphing a function and noticed that at certain points, the graph reaches invisible lines the graph will never cross. Explain to John what the two types of invisible lines are and how to predict them. You may create your own example to aid in your reasoning. Use complete sentences.

OpenStudy (anonymous):

These would be horizontal and vertical asymptotes, do you know how to find these?

OpenStudy (redheadangel):

The "invisible lines" are asymptotes.

OpenStudy (redheadangel):

like for example y=xx−1 would have a vertical asymptote at x=1

OpenStudy (redheadangel):

horizontal ones are what the function approaches as x goes to ∞ or −∞ for example y=xx−1 would have a horizontal asymptote at y=1

OpenStudy (redheadangel):

if you have a function that is one polynomial over another, a "rational function" then the vertical asymptote is where the denominator is equal to zero for example if i had y=x+3x−2 and set x−2=0 i get x=2 and the vertical line x=2 is the vertical asymptote